OpenStudy (anonymous):

medals for answer wat do you get when you put 7 and 7 together

OpenStudy (anonymous):

77

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

define this "put" function. i'm not familiar with it

OpenStudy (anonymous):

well just put them together its very easy

OpenStudy (anonymous):

you can get a multitude of answers. define "put"

OpenStudy (anonymous):

you guys have to guess come come now if i told you the answer will be given come on plz

OpenStudy (anonymous):

7+7=14 7-7=0 7*7=49 7/7=1

OpenStudy (anonymous):

it has nothing to do with math

OpenStudy (anonymous):

its a dame triganel man geeezzz